# Euclid Quick Data Release (Q1)

### v2

@@ 2.3 Training Objective
For training, we use $\times64$ pixel VIS cutouts for each of the selected sources that will be discussed in Sect. 4.1. We use the training pipeline described in Dhariwal & Nichol (2021) [^1] Available at https://github.com/openai/guided-diffusion and adopt a hybrid loss function that combines a simple mean squared error (MSE) term with the full variational lower bound (VLB).
The MSE objective, proposed by Ho et al. (2020), is an MSE loss between the true and predicted Gaussian noise, $\epsilon$
@@
where $x_{0}$ is an original image, $t$ is a uniformly sampled timestep, and $\epsilon\sim N(0,I)$ is the true noise. For our implementation $\bar{\alpha}_{t}$ is described below, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
@@
Since $L_{VLB}$ explicitly depends on the parameters of $\Sigma_{\theta}$, it can be used to train them. We adopt the hybrid objective
$L_{\text{hybrid}}=L_{\text{MSE}}+\lambda L_{\text{VLB}}\;.$ (8)
We use a small weighting constant ($\lambda=0.001$) to prevent the $L_{\mathrm{VLB}}$ term from overpowering the $L_{\mathrm{MSE}}$ term, which remains the primary driver of sample quality. This hybrid loss allows the model to learn an optimal reverse process variance, which can improve model likelihoods and enable more efficient sampling. A key challenge with astronomical data is its high dynamic range. A standard MSE-based loss is not scale-invariant and will be dominated by the few brightest pixels or sources in a training batch. This can lead to poor performance on the much larger population of fainter objects. To address this, we introduce a normalised version of the MSE loss term. For each image $x_{i}$, we divide its error by its respective maximum pixel value, denoted as $\max(x_{i})$. This normalisation ensures that the relative error is penalised, allowing for more equitable training across both bright and faint sources. The final loss used for our primary model is thus
$L_{\text{normalised}}=\frac{L_{\text{MSE}}}{\max(x_{i})}+\lambda L_{\text{VLB}}\;.$ (9)
Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss. The training curves of this scale-invariant loss compared with the original loss used in Nichol & Dhariwal (2021) is shown in Fig. 18. By incorporating this normalised loss, we enable more equitable training across the full range of galaxy brightnesses, rather than being dominated by the brightest objects which occurs when using the default loss.
@@ 2.4 Noise scheduling
The total number of timesteps, $T$, is a critical hyperparameter in the diffusion process. A larger $T$ corresponds to smaller, more refined steps in both the forward (noising) and reverse (denoising) processes. While this can, in principle, allow the model to learn a more accurate representation of the data distribution, it comes with a significant trade-off. The computational cost of inference (i.e., generating a sample) is directly proportional to $T$, as the reverse process must be iterated $T$ times. The cost of a single training step, however, is not directly dependent on $T$, because at each step a random timestep $t$ is sampled and the loss is evaluated only for that level of noise.
Figure 2: The linear schedule from the original diffusion implementation causes the parameters to converge early in the timesteps, resulting in training images becoming pure noise too soon and leading to suboptimal performance. The switch to the cosine-beta schedule adds noise at a much slower rate, prioritising smaller updates in the early stages leading to more unique noised images throughout training.
In the original DDPM paper, Ho et al. (2020) made use of a linear schedule for the amount of noise added to the image at each $t$. Although the produced images competed with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low-resolution images and proposed using a cosine-beta noise schedule as a replacement. By stretching the noise levels to add less noise in early iterations and more in the latter, the process overcame the issue of images becoming too noisy too quickly. Figure 2 shows how the scheduler impacts the $\beta$ parameter used in Eq. 2. Equation 10 shows how $\beta_{t}$ is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
@@ 2.5 Conditioning the model
Using the pipeline of DDPMs allows us to impose conditioning on our input space. Throughout training, the model is already conditioned on the timestep $t$. This conditioning is crucial, as it informs the neural network of the noise level present in the input image $x_{t}$. The network learns a single function, $\epsilon_{\theta}(x_{t},t)$, that is trained to predict the noise component $\epsilon$ from a noised image $x_{t}$ at any given timestep $t$. The characteristics of the noise to be removed differ substantially between a timestep close to $T$ (where the image is almost pure noise) and a timestep close to 0 (where the image is almost fully reconstructed). Providing $t$ as an input allows the single network to adapt its predictions accordingly.
@@
However, we are not limited to only conditioning by timestep; often, multiple parameters are provided alongside each image. Typically, this is to allow ‘class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al. 2021; Zhang et al. 2023). By providing an embedding space during training, the model can learn not only the dynamics of the noise but also the inherent differences between noise and non-noise pixel values for each of the class representations. This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class. Providing rich representations to the model has led to the impressive text-to-image capabilities of popular diffusion-based implementations, such as Dall-E (Ramesh et al. 2022) or Imagen (Saharia et al. 2022).
@@
This type of conditioning allows the user to direct the image generation to a particular class or style of image. However, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt will likely produce an image very different from the first, even if the generated image belongs to the requested class. Even with the possibility of detailed prompts and refined classes, which can guide the generation towards a narrower or more specific region of the learned space, it may be difficult to know exactly what the output image will look like. This is due to the one-to-many relationship between classes/prompts and their subsequent generated outputs.
@@
However, for our task, we require a different form of conditioning. Instead of generating an entire image based on a class label, we need to reconstruct a specific, masked region of an existing image, conditioned on the surrounding, unmasked pixels. This task is known as inpainting (Pathak et al. 2016; Yeh et al. 2017; Yu et al. 2018; Wang et al. 2018). To achieve this, we do not train a new, specialised conditional model. Instead, we employ a ‘plug-and-play’ (PnP) approach, where conditioning is introduced during the inference phase of an unconditionally trained generative model (Nguyen et al. 2017; Graikos et al. 2022). Our diffusion model, trained on a vast sample of galaxies, serves as a powerful, non-parametric prior for galaxy morphology. The PnP approach allows us to apply this prior to the inpainting task with great flexibility, as the model does not need to be retrained for different images or mask geometries. This is particularly advantageous for large-scale surveys like Euclid. Recent works in astrophysics have successfully employed similar PnP or posterior sampling strategies for complex inference tasks (Feng et al. 2023; Dia et al. 2025).
@@ 2.6 Repainting
For this work, we use the Repaint algorithm (Lugmayr et al. 2022), a PnP method specifically designed for diffusion models. The core challenge in diffusion-based inpainting is to ensure semantic and structural consistency between the known (unmasked) and unknown (masked) regions. A naive approach of simply replacing the masked region with noise and then running the reverse process would fail, as the denoising network would alter the known, unmasked pixels, corrupting the conditioning information.
Repaint overcomes this by iteratively enforcing the data constraint at each step $t$ of the reverse diffusion process. Given an image $x_{0}$ and a binary mask $M$ (where $1$ indicates the region to be inpainted and 0 indicates the known region), the algorithm proceeds as follows. At each timestep $t$ of the reverse process, starting from a noisy sample $x_{t}$:
@@
    Sample the known region: The original image, $x_{0}$, is noised to the appropriate level for timestep $t$ using the forward process $q(x_{t}|x_{0})$, resulting in a sample of the known part, $x^{\mathrm{known}}_{t}$.
@@
    Sample the unknown region: The output image from the previous, slightly more noisy step, $x_{t+1}$, is denoised for one step using the unconditional diffusion model to get a sample of the unknown part, $x^{\mathrm{unknown}}_{t}$.
@@
This procedure, illustrated in Fig. 1, ensures that the generated region is coherently conditioned on the surrounding galaxy structure at every stage of the generation. The ‘denoise’ operation in this context refers to the application of one step of the learned reverse diffusion process, $p_{\theta}(x_{t-1}|x_{t})$, which involves using the trained neural network $\epsilon_{\theta}$ to predict the noise in $x_{t}$ and deriving an estimate for $x_{t-1}$. By denoising an image composed of both fixed and generated pixels – each matched to the appropriate noise level $t$ – we ensure that the generated region is properly conditioned on the information from the rest of the galaxy, resulting in a final image where all pixels are consistent and coherent.
While alternative conditioning methods based on guiding the reverse process with a likelihood score function exist (Graikos et al. 2022), we adopt the Repaint algorithm for its straightforward implementation and proven effectiveness in enforcing data consistency in a plug-and-play setting.
@@ 3 Training and inference pipeline
The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting. Training the diffusion model uses the whole image and learns to predict what noise to remove at each timestep, allowing for generating novel examples from the trained distribution.
@@ 3.1 Mask creation
Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time. Each mask is positioned so that the brightest pixel in the centre of the source becomes the centre of the mask. Although each image is centred on the source, various interactions and dynamics allow suspected AGN to be offset from the centre. To allow for some deviation, a $\times9$ pixel window around the centre of the image is used to detect the brightest pixel. This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake. The masks for each image are assigned automatically throughout the data-loading pipeline.
@@ 3.2 Reconstruction rescaling
Unlike some image processing pipelines that rescale image data to a limited range, such as when working with JPG images which are binned to the standard 0–255 pixel scale range and converted to floats between 0–1 (Walmsley et al. 2020; Euclid Collaboration: Aussel et al. 2024), our approach utilises the raw values direct from the sensor as input to our network. As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range. Although this did not prevent the model from effectively learning the distribution, it does require a rescaling of the generated output as the pixels are mostly between -1 and ++1 to improve training stability.
@@
To allow the inpainted region to be meaningfully compared (e.g., via MSE) to the original pixel values in that masked region, a post-inference rescaling must be applied. The vast majority of pixels in the image (the unmasked region, which constitutes 99.5% of a $\times64$ image when a $\times5$ central mask is used) correspond to known pixel values from the original astronomical image. We leverage these known, unmasked pixels to perform a linear rescaling of the diffusion model’s initial output (which is in a normalised range) back to the original image’s dynamic range. This procedure ensures consistency in the unmasked regions and appropriately scales the newly inpainted pixels in the masked region. The following scaling can be applied instantly on output of the repaint procedure using each image’s respective values to allow the non-masked pixels of the input and output images to be consistent:
@@ 4.1 Sample selection
Covering 63.1 deg2, the Euclid Q1 Data Release (Euclid Quick Release Q1 2025) contains just under 30 million catalogued sources (Euclid Collaboration: Romelli et al. 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F). In this work, we use exclusively the VIS images (Euclid $I_{\scriptscriptstyle\rm E}$ band) because we aim to identify AGN candidates on a single image. However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies. The following criteria are applied to the 30 million source catalogue of Q1:
@@
  • the source must have a $I_{\scriptscriptstyle\rm E}$ magnitude brighter than $22.5$ (23.5 - 2.5log(flux_detection_total_euclid) << 22.5);
@@
  • the signal-to-noise ratio of the total flux measurement must be sufficiently high ($\frac{\texttt{flux\_detection\_total\_euclid}}{\texttt{fluxerr\_detection\_total\_euclid}}$ >> 15); and
  • if the signal-to-noise criterion is not met, then the segmentation area of the source must be very large (segmentation_area_euclid >> 1200), as defined in Euclid Collaboration: Walmsley et al. (2025).
@@
After this sample selection, we are left with 1 142 606 sources across the three Q1 fields. VIS image cutouts ($\times64$ pixels) were created for each source. Problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed. This provided us with images and catalogue data for 1 060 864 sources. We then apply a 80 : 20 train and test split. Even though the inpainting and its results are a separate pipeline to the diffusion training, all of the decision boundaries created are based solely on the training set to prevent any potential data leakage. When showing classifier performance, all results are from the unseen 20% of the data. This prevents any potential memorisation of the training images from impacting the results. However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.

### v3

@@ 2.3 Training objective
For training, we used $\times64$ pixel VIS cutouts for each of the selected sources that are discussed in Sect. 4.1. We used the training pipeline described in Dhariwal & Nichol (2021) [^2] Available at https://github.com/openai/guided-diffusion and adopted a hybrid loss function that combines a simple mean squared error (MSE) term with the full variational lower bound (VLB).
The MSE objective, proposed by Ho et al. (2020), is an MSE loss between the true and predicted Gaussian noise, $\epsilon$,
@@
where $x_{0}$ is an original image, $t$ is a uniformly sampled timestep, and $\epsilon\sim N(0,I)$ is the true noise. For our implementation $\bar{\alpha}_{t}$ is described below (the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation):
@@
Since $L_{\mathrm{VLB}}$ explicitly depends on the parameters of $\Sigma_{\theta}$, it can be used to train them. We adopted the hybrid objective
$L_{\mathrm{hybrid}}=L_{\mathrm{MSE}}+\lambda L_{\mathrm{VLB}}\;.$ (8)
We used a small weighting constant ($\lambda=0.001$) to prevent the $L_{\mathrm{VLB}}$ term from overpowering the $L_{\mathrm{MSE}}$ term, which remains the primary driver of sample quality. This hybrid loss allows the model to learn an optimal reverse process variance, which can improve model likelihoods and enable more efficient sampling. A key challenge with astronomical data is its high dynamic range. A standard MSE-based loss is not scale-invariant and will be dominated by the few brightest pixels or sources in a training batch. This can lead to poor performance on the much larger population of fainter objects. To address this, we introduced a normalised version of the MSE loss term. For each image, $x_{i}$, we divided its error by its respective maximum pixel value, denoted as $\max(x_{i})$. This normalisation ensures that the relative error is penalised, allowing for more equitable training across both bright and faint sources. The final loss used for our primary model is thus
$L_{\mathrm{normalised}}=\frac{L_{\mathrm{MSE}}}{\max(x_{i})}+\lambda L_{\mathrm{VLB}}\;.$ (9)
Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss. The training curves of this scale-invariant loss compared with the original loss used in Nichol & Dhariwal (2021) are shown in Fig. 18. By incorporating this normalised loss, we enable more equitable training across the full range of galaxy brightnesses, rather than being dominated by the brightest objects, which occurs when using the default loss.
@@ 2.4 Noise scheduling
The total number of timesteps, $T$, is a critical hyperparameter in the diffusion process. A larger $T$ corresponds to smaller, more refined steps in both the forward (noising) and reverse (denoising) processes. While this can, in principle, allow the model to learn a more accurate representation of the data distribution, it comes with a significant trade-off. The computational cost of inference (i.e. generating a sample) is directly proportional to $T$, as the reverse process must be iterated $T$ times. The cost of a single training step, however, is not directly dependent on $T$, because at each step a random timestep, $t$, is sampled and the loss is evaluated only for that level of noise.
Figure 2: Linear schedule from the original diffusion implementation, which causes the parameters to converge early in the timesteps, resulting in training images becoming pure noise too soon and leading to suboptimal performance. The switch to the cosine-beta schedule adds noise at a much slower rate, prioritising smaller updates in the early stages, leading to more unique noised images throughout training.
In the original DDPM paper, Ho et al. (2020) made use of a linear schedule for the amount of noise added to the image at each $t$. Although the produced images competed with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low-resolution images and proposed using a cosine-beta noise schedule as a replacement. By stretching the noise levels to add less noise in early iterations and more in the latter, the process overcame the issue of images becoming too noisy too quickly. Figure 2 shows how the scheduler impacts the $\beta$ parameter used in Eq. 2. Equation 10 shows how $\beta_{t}$ is calculated (the reader is directed to Nichol & Dhariwal 2021 for specifics on its derivation):
@@ 2.5 Conditioning the model
Using the pipeline of DDPMs allows us to impose conditioning on our input space. Throughout training, the model is already conditioned on the timestep, $t$. This conditioning is crucial, as it informs the neural network of the noise level present in the input image, $x_{t}$. The network learns a single function, $\epsilon_{\theta}(x_{t},t)$, which is trained to predict the noise component, $\epsilon$, from a noised image, $x_{t}$, at any given timestep, $t$. The characteristics of the noise to be removed differ substantially between a timestep close to $T$ (where the image is almost pure noise) and a timestep close to 0 (where the image is almost fully reconstructed). Providing $t$ as an input allows the single network to adapt its predictions accordingly.
@@
However, we are not limited to only conditioning by timestep; often, multiple parameters are provided alongside each image. Typically, this is to allow ‘class-conditioning’, whereby users can tune the expected output of the diffusion model to a specific class or category (Radford et al., 2021; Zhang et al., 2023). By providing an embedding space during training, the model can learn not only the dynamics of the noise but also the inherent differences between noise and non-noise pixel values for each of the class representations. This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class. Providing rich representations to the model has led to the impressive text-to-image capabilities of popular diffusion-based implementations, such as Dall-E (Ramesh et al., 2022) or Imagen (Saharia et al., 2022).
@@
This type of conditioning allows the user to direct the image generation to a particular class or style of image. However, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt will likely produce an image very different from the first, even if the generated image belongs to the requested class. Even with the possibility of detailed prompts and refined classes, which can guide the generation towards a narrower or more specific region of the learned space, it may be difficult to know exactly what the output image will look like. This is due to the one-to-many relationship between classes or prompts and their subsequent generated outputs.
@@
However, for our task, we require a different form of conditioning. Instead of generating an entire image based on a class label, we need to reconstruct a specific, masked region of an existing image, conditioned on the surrounding, unmasked pixels. This task is known as inpainting (Pathak et al., 2016; Yeh et al., 2017; Yu et al., 2018; Wang et al., 2018). To achieve this, we do not train a new, specialised conditional model. Instead, we employ a ‘plug-and-play’ (PnP) approach, whereby conditioning is introduced during the inference phase of an unconditionally trained generative model (Nguyen et al., 2017; Graikos et al., 2022). Our diffusion model, trained on a vast sample of galaxies, serves as a powerful, non-parametric prior for galaxy morphology. The PnP approach allows us to apply this prior to the inpainting task with great flexibility, as the model does not need to be retrained for different images or mask geometries. This is particularly advantageous for large-scale surveys like Euclid. Recent works in astrophysics have successfully employed similar PnP or posterior sampling strategies for complex inference tasks (Feng et al., 2023; Dia et al., 2025).
@@ 2.6 Repainting
For this work, we used the Repaint algorithm (Lugmayr et al., 2022), a PnP method specifically designed for diffusion models. The core challenge in diffusion-based inpainting is to ensure semantic and structural consistency between the known (unmasked) and unknown (masked) regions. A naive approach of simply replacing the masked region with noise and then running the reverse process would fail, as the denoising network would alter the known, unmasked pixels, corrupting the conditioning information.
Repaint overcomes this by iteratively enforcing the data constraint at each step, $t$, of the reverse diffusion process. Given an image, $x_{0}$, and a binary mask, $M$ (where $1$ indicates the region to be inpainted and 0 indicates the known region), the algorithm proceeds as follows. At each timestep, $t$, of the reverse process, starting from a noisy sample, $x_{t}$:
@@
    Sample the known region: The original image, $x_{0}$, is noised to the appropriate level for timestep $t$ using the forward process, $q(x_{t}|x_{0})$, resulting in a sample of the known part, $x^{\mathrm{known}}_{t}$.
@@
    Sample the unknown region: The output image from the previous, slightly noisier step, $x_{t+1}$, is denoised for one step using the unconditional diffusion model to get a sample of the unknown part, $x^{\mathrm{unknown}}_{t}$.
@@
This procedure, illustrated in Fig. 1, ensures that the generated region is coherently conditioned on the surrounding galaxy structure at every stage of the generation. The ‘denoise’ operation in this context refers to the application of one step of the learned reverse diffusion process, $p_{\theta}(x_{t-1}|x_{t})$, which involves using the trained neural network, $\epsilon_{\theta}$, to predict the noise in $x_{t}$ and deriving an estimate for $x_{t-1}$. By denoising an image composed of both fixed and generated pixels – each matched to the appropriate noise level $t$ – we ensure that the generated region is properly conditioned on the information from the rest of the galaxy, resulting in a final image in which all pixels are consistent and coherent.
While alternative conditioning methods based on guiding the reverse process with a likelihood score function exist (Graikos et al., 2022), we adopted the Repaint algorithm for its straightforward implementation and proven effectiveness in enforcing data consistency in a PnP setting.
@@ 3 Training and inference pipeline
The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting. Training the diffusion model uses the whole image and learns to predict what noise to remove at each timestep, allowing for novel examples to be generated from the trained distribution.
@@ 3.1 Mask creation
Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time. Each mask is positioned so that the brightest pixel in the centre of the source becomes the centre of the mask. Although each image is centred on the source, various interactions and dynamics allow suspected AGN to be offset from the centre. To allow for some deviation, a $\times9$ pixel window around the centre of the image is used to detect the brightest pixel. This offset allows for the correct masking of less symmetric sources, whilst limiting the likelihood of masking an adjacent source by mistake. The masks for each image are assigned automatically throughout the data-loading pipeline.
@@ 3.2 Reconstruction rescaling
Unlike some image processing pipelines that rescale image data to a limited range, such as when working with JPG images that are binned to the standard 0–255 pixel scale range and converted to floats between 0–1 (Walmsley et al., 2020; Euclid Collaboration: Aussel et al., 2024), our approach utilises the raw values direct from the sensor as input to our network. As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
One potential issue with this set-up is that the libraries used for training are optimised for images within the standard pixel range. Although this did not prevent the model from effectively learning the distribution, it does require a rescaling of the generated output as the pixels are mostly between -1 and ++1 to improve training stability.
@@
To allow the inpainted region to be meaningfully compared (e.g. via MSE) to the original pixel values in that masked region, a post-inference rescaling must be applied. The vast majority of pixels in the image (the unmasked region, which constitutes 99.5% of a $\times64$ image when a $\times5$ central mask is used) correspond to known pixel values from the original astronomical image. We leveraged these known, unmasked pixels to perform a linear rescaling of the diffusion model’s initial output (which is in a normalised range) back to the original image’s dynamic range. This procedure ensures consistency in the unmasked regions and appropriately scales the newly inpainted pixels in the masked region. The following scaling can be applied instantly on output of the repaint procedure using each image’s respective values to allow the non-masked pixels of the input and output images to be consistent:
@@ 4.1 Sample selection
Covering 63.1 deg2, the Euclid Q1 Data Release (Euclid Quick Release Q1, 2025) contains just under 30 million catalogued sources (Euclid Collaboration: Romelli et al., 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F). In this work, we use exclusively the VIS images (Euclid $I_{\scriptscriptstyle\rm E}$ band) because we aim to identify AGN candidates on a single image. However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies. The following criteria are applied to the 30 million source catalogue of Q1:
@@
  • the source must have a $I_{\scriptscriptstyle\rm E}$ magnitude brighter than $22.5$ [23.5 - 2.5log(flux_detection_total_euclid) << 22.5];
@@
  • the S/N of the total flux measurement must be sufficiently high ($\frac{\texttt{flux\_detection\_total\_euclid}}{\texttt{fluxerr\_detection\_total\_euclid}}$ >> 15); and
  • if the S/N criterion is not met, then the segmentation area of the source must be very large (segmentation_area_euclid >> 1200), as defined in Euclid Collaboration: Walmsley et al. (2025).
@@
After this sample selection, we are left with 1 142 606 sources across the three Q1 fields. VIS image cutouts ($\times64$ pixels) were created for each source. Problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed. This provided us with images and catalogue data for 1 060 864 sources. We then applied a 80 : 20 train and test split. Even though the inpainting and its results are a separate pipeline to the diffusion training, all of the decision boundaries created are based solely on the training set to prevent any potential data leakage. When showing classifier performance, all results are from the unseen 20% of the data. This prevents any potential memorisation of the training images from impacting the results. However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.